40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5418     Entry time: Thu Sep 15 16:45:59 2011
Author: Paul 
Type: Update 
Category: SUS 
Subject: ITMY and SRM Oplev status 

Today I worked on getting the ITMY and SRM oplevs back in working order. I aligned the SRM path back onto the QPD. I put excitations on the ITMY and SRM in pitch and yaw and observed the beam at the QPDs to check for clipping. They looked clean from clipping.

 
Measurements of the beam power at various points:
 
Straight after the laser - 7.54mW
After the BS in the SRM path - 1.59mW
After the BS in the ITMY path - 3.24mW
Incident on the SRM QPD - 0.03mW
Incident on the ITMY QPD - 0.25mW
 
Counts registered from the QPD sum channels:
 
SRM QPD SUM dark count - 1140
SRM QPD SUM bright count - 3250
 
ITMY QPD SUM dark count - 150
ITMY QDP SUM bright count - 12680
 
The power incident on the SRM QPD seems very low with respect to the ITMY QPD. Is the SRM mirror coating not very reflective for the He-Ne laser?There are some back reflections from lenses, which we should be careful of to avoid scattering.
ELOG V3.1.3-